OpenStudy (anonymous):

FAN & MEDAL A part-time house cleaner made $9258.13 last year. If he claimed himself as an exemption for $3650 and had a $5700 standard deduction, what was his taxable income last year? A. $91.87 B. $3558.13 C. $5608.13 D. $0

OpenStudy (anonymous):

9258.13 - (3650 + 5700) = 9258.13 - 9350 = - 91.87 If I did this correctly, he should not have any taxable income :)

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

I hope it is correct.....I think it is :)

OpenStudy (anonymous):

It is, thank you!

OpenStudy (anonymous):

thats incorrect it was 0